# Convergence of a Finite Volume Scheme for Compactly Heterogeneous Scalar Conservation Laws

### v2

@@ 3.1 Discontinuous Flux
In this section, we recall and adapt for our use standard results regarding conservation laws with discontinuous flux. For the purpose of this section, let us fix $f_{l},f_{r}\in\mathbf{C}^{1}(\mathbb{R},\mathbb{R})$ two convex functions satisfying (CVX). Consider:
@@
Since the works of [1, 2, 3, 4, 13, 19, 24, 26, 37], it is now well known that additionally to the conservation of mass (Rankine-Hugoniot condition)
[…]
an entropy criterion must be imposed at the interface to select one solution. This choice is often guided by physical consideration. Among the most common criteria, we can cite: the minimal jump condition [19, 23, 22], the vanishing viscosity criterion [5], or the flux maximization [3].
If $\Phi_{l}$, resp. $\Phi_{r}$, denotes the Kruzhkov entropy flux associated with $f_{l}$, resp. with $f_{r}$, then in this section, define $\Phi=\Phi(x,u)$ as the Kruzhkov entropy flux associated with $F$:
@@
For the study of (3.1), we follow [4], where the traces of the solution at the interface $\{x=0\}$ are explicitly treated. It will be useful to have a name for the critical points of $f_{l,r}$, say $\alpha_{l,r}$. Notice that for all $y\in\mathopen{]}\min f_{r},+\infty\mathclose{[}$, the equation $f_{r}(u)=y$ admits exactly two solutions, $S_{r}^{-}(y)<\alpha_{r}<S_{r}^{+}(y)$. When $y=\min f_{r}$, then $S_{r}^{-}(y)=S_{r}^{+}(y)=\alpha_{r}$. This motivates the following definition.
@@
The admissibility germ for (3.1) is the subset $\mathcal{G}$ defined as the union of:
$\mathcal{G}_{1}$ :-$\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}\geq\alpha_{l},\;k_{r}=S_{r}^{+}(f_{l}(k_{l}))\right\}:$ (3.2)
$\mathcal{G}_{2}$ :-$\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}\leq\alpha_{l},\;k_{r}=S_{r}^{-}(f_{l}(k_{l}))\right\}:$
$\mathcal{G}_{3}$ :-$\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}>\alpha_{l},\;k_{r}=S_{r}^{-}(f_{l}(k_{l}))\right\}.:$
The germ contains all the possible traces along $\{x=0\}$ of the solutions to (3.1). Notice that by construction, any couple in the germ satisfies the Rankine-Hugoniot condition. Conversely, some couples verifying the Rankine-Hugoniot condition have been excluded from the germ, more precisely the ones belonging to
@@
(ii) Dissipative inequality:
$\forall(u_{l},u_{r}),(k_{l},k_{r})\in\mathcal{G},\quad\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0.$ (3.3)
@@
$\forall(k_{l},k_{r})\in\mathcal{G},\quad\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0\implies(u_{l},u_{r})\in\mathcal{G}.$ (3.4)
Proof.  Point (i) is clear, by construction, while (ii) follows from a straightforward case by case study. We now prove (iii) by way of contradiction: let $(u_{l},u_{r})\in\mathbb{R}^{2}$ such that $f_{l}(u_{l})=f_{r}(u_{r})$ and assume that
@@
From the Rankine-Hugoniot condition, we deduce that $u_{r}=S_{r}^{+}(f_{l}(u_{l}))$, with $u_{l}<\alpha_{l}$. Take $k_{r}=S_{r}^{-}(f_{(}k_{l}))$ with $k_{l}\in\mathopen{]}u_{l},\alpha_{l}\mathclose{[}$. Clearly, $(k_{l},k_{r})\in\mathcal{G}_{2}\subset\mathcal{G}$, therefore by assumption, $\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0$. But the choice of $k_{l}$ raises the contradiction:
@@
The theory of [4] proposes an abstract framework for the study of (3.1). On the other hand, the authors of [3, 2] give a formula for the flux at the interface for Riemann problems of (3.1). In the following proposition, we link the two points of view. Let us adopt the notations:
@@
$\mathcal{R}(u_{l},u_{r})\coloneq|F_{\mathrm{int}}(u_{l},u_{r})-f_{l}(u_{l})|+|F_{\mathrm{int}}(u_{l},u_{r})-f_{r}(u_{r})|.$ (3.7)
@@
(ii) For clarity, set $\mu\coloneq\max\{\min f_{l},\min f_{r}\}$ and notice that
$\mathcal{R}(u_{l},u_{r})=\left\{\begin{aligned} |f_{l}(u_{l})-\mu|+|f_{r}(u_{r})-\mu|&\quad\text{if}\;u_{l}<\alpha_{l}\;\text{and}\;u_{r}\geq\alpha_{r}\\ |f_{l}(u_{l})-f_{r}(u_{r})|&\quad\text{otherwise.}\end{aligned}\right.$
Let us call $q\coloneq\Phi_{r}(u_{r},k_{r})-\Phi_{l}(u_{l},k_{l})$. We proceed with a case by case study. We use the fact that $f_{l,r}$ is decreasing on $\mathopen{]}-\infty,\alpha_{l,r}]$ and increasing on $[\alpha_{l,r},+\infty\mathclose{[}$.
@@
Case 3: $u_{l}>k_{l}$ and $u_{r}\leq k_{r}$. Then $q=(f_{r}(k_{r})-f_{r}(u_{r}))+(f_{l}(k_{l})-f_{l}(u_{l}))$. Notice that because of the germ structure, the case $k_{l}<u_{l}<\alpha_{l}$ and $\alpha_{r}\leq u_{r}\leq k_{r}$ cannot happen. Therefore, in Case 3, $\mathcal{R}(u_{l},u_{r})=|f_{l}(u_{l})-f_{r}(u_{r})|$. We have
$\bullet$ $k_{l}<u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}\leq k_{r}\leq\alpha_{r}$
$\implies q\leq f_{l}(k_{l})-f_{l}(u_{l})=f_{r}(u_{r})-f_{l}(u_{l})+(f_{r}(k_{r})-f_{r}(u_{r}))\leq f_{r}(u_{r})-f_{l}(u_{l})\leq\mathcal{R}(u_{l},u_{r})$
$k_{l}<u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}\leq\alpha_{r}<k_{r}\;\text{does not happen}$
$\bullet$ $k_{l}<u_{l}\leq\alpha_{l}\;\text{and}\;\alpha_{r}<u_{r}\leq k_{r}\;\text{does not happen}$
$\bullet$ $\alpha_{l}<k_{l}<u_{l}\;\text{and}\;u_{r}\leq k_{r}\leq\alpha_{r}\implies q\leq 0$
$k_{l}<\alpha_{l}<u_{l}\;\text{and}\;u_{r}\leq k_{r}\leq\alpha_{r}$
$\implies q\leq f_{l}(k_{l})-f_{l}(u_{l})=f_{r}(u_{r})-f_{l}(u_{l})+(f_{r}(k_{r})-f_{r}(u_{r}))\leq f_{r}(u_{r})-f_{l}(u_{l})\leq\mathcal{R}(u_{l},u_{r})$
$\alpha_{l}<k_{l}<u_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\leq k_{r}$
$\implies q\leq f_{r}(k_{r})-f_{r}(u_{r})=f_{l}(u_{l})-f_{r}(u_{r})+(f_{l}(k_{l})-f_{l}(u_{l}))\leq f_{l}(u_{l})-f_{r}(u_{r})\leq\mathcal{R}(u_{l},u_{r})$
$\alpha_{l}<k_{l}<u_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\leq k_{r}\;\text{does not happen}$
$\bullet$ $\alpha_{l}<k_{l}<u_{l}\;\text{and}\;\alpha_{r}<u_{r}\leq k_{r}$
$\implies q\leq f_{r}(k_{r})-f_{r}(u_{r})=f_{l}(u_{l})-f_{r}(u_{r})+(f_{l}(k_{l})-f_{l}(u_{l}))\leq f_{l}(u_{l})-f_{r}(u_{r})\leq\mathcal{R}(u_{l},u_{r})$
$k_{l}<\alpha_{l}<u_{l}\;\text{and}\;\alpha_{r}<u_{r}\leq k_{r}\;\text{does not happen.}$.
When we say “does not happen”, we mean that germ geometry prevents the situation from occurring.
Case 4: $u_{l}\leq k_{l}$ and $u_{r}>k_{r}$. Similar to Case 3.   $\square$
@@
$\mathcal{R}(k,k)=\left\{\begin{aligned} |\max\{\min f_{l},\min f_{r}\}-f_{l}(k)|+|\max\{\min f_{l},\min f_{r}\}-f_{r}(k)|&\quad\text{if}\;\alpha_{r}\leq k<\alpha_{l}\\ |f_{l}(k)-f_{r}(k)|&\quad\text{otherwise.}\end{aligned}\right.$
We are now in position to properly give the definition of solutions for (3.1).
@@
Define $\mathcal{G}$ as (3.2) and $\mathcal{R}$ as (3.7). Let $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$ and $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$. Then the following statements are equivalent.
1. For all test functions $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ and for all $(k_{l},k_{r})\in\mathbb{R}^{2}$,
$\int_{0}^{+\infty}\int_{\mathbb{R}}|u-\kappa(x)|\partial_{t}\varphi+\Phi(x,u,\kappa(x))\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}|u_{o}(x)-\kappa(x)|\varphi(0,x)\mathinner{\mathrm{d}{x}}$
@@
2.(i) For all test functions $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R}^{*},\mathbb{R}^{+})$ and for all $(k_{l},k_{r})\in\mathbb{R}^{2}$,
$\int_{0}^{+\infty}\int_{\mathbb{R}}|u-\kappa(x)|\partial_{t}\varphi+\Phi(x,u,\kappa(x))\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}|u_{o}(x)-\kappa(x)|\varphi(0,x)\mathinner{\mathrm{d}{x}}\geq 0.$ (3.9)
2.(ii) For a.e. $t\in\mathopen{]}0,+\infty\mathclose{[}$, $(u(t,0-),u(t,0+))\in\mathcal{G}$.
When one of these statements holds, we say that $u$ is an entropy solution to (3.1) with initial datum $u_{o}$.
Proof.  See the proof of [4, Theorem 3.18].   $\square$
Remark 3.2.
In both (3.4) and (3.9), $\kappa$ denotes the piecewise constant function
$\kappa(x)=\left\{\begin{aligned} k_{l}&\quad\text{if}\;x\leq 0\\ k_{r}&\quad\text{if}\;x>0.\end{aligned}\right.$
In the second definition, the traces of the solution along $\{x=0\}$ explicitly appear. The condition
@@
is to be understood in the sense of strong traces. An important fact we stress is that it is not restrictive to assume that general entropy solutions, that is to say bounded functions verifying (3.9), admit strong traces. Usually, it is ensured provided a nondegeneracy assumption on the flux function. Here, since $f_{l}$ and $f_{r}$ are genuinely nonlinear in the sense that
$\forall s\in\mathbb{R},\quad\text{meas}\biggl{(}\left\{p\in\mathbb{R}\;:\;f_{l,r}^{\prime}(p)=s\right\}\biggr{)}=0,$
existence of strong traces is well known since [38], see also [33].
@@
    Let $u_{o},v_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. We denote by $u$, resp. $v$, an entropy solution to (3.1) with initial datum $u_{o}$, resp. $v_{o}$. Then, there exists $L>0$ such that for all $R>0$ and for all $t>0$,
    $\int_{|x|\leq R}|u(t,x)-v(t,x)|\mathinner{\mathrm{d}{x}}$ $\leq\int_{|x|\leq R+Lt}|u_{o}(x)-v_{o}(x)|\mathinner{\mathrm{d}{x}}$ (3.10)
    $\int_{|x|\leq R}(u(t,x)-v(t,x))^{+}\mathinner{\mathrm{d}{x}}$ $\leq\int_{|x|\leq R+Lt}(u_{o}(x)-v_{o}(x))^{+}\mathinner{\mathrm{d}{x}}.$
Proof.  The proofs of the stability bounds can be found in [4, Theorems 3.11-3.19]. Regarding the existence, one can build a simple finite volume scheme and prove its convergence, see for instance [2]. The flux to put at the interface $\{x=0\}$ in the scheme is precisely $F_{\mathrm{int}}$.   $\square$
@@
$F_{\mathrm{int}}(u_{l},u_{r})=\left\{\begin{array}[]{ccl}\max\{f_{l}(\alpha_{l}),f_{r}(u_{r})\}&\text{if}&u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\quad\text{(I)}\\ \max\{\min f_{l},\min f_{r}\}&\text{if}&u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}>\alpha_{r}\quad\text{(II)}\\ \max\{f_{l}(u_{l}),f_{r}(u_{r})\}&\text{if}&u_{l}>\alpha_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\quad\text{(III)}\\ \max\{f_{l}(u_{l}),\min f_{r}\}&\text{if}&u_{l}>\alpha_{l}\;\text{and}\;u_{r}>\alpha_{r}\quad\text{(IV)}\end{array}\right.$ (3.11)
The value of $F_{\mathrm{int}}(u_{l},u_{r})$ imposes the value of one of the traces at the interface. For instance, if $F_{\mathrm{int}}(u_{l},u_{r})=f_{r}(u_{r})$, then the right trace of the solution along $\{x=0\}$ is equal to $u_{r}$. To determine the left trace, one solves $f_{l}(\gamma)=f_{r}(u_{r})$, $\gamma\in\mathbb{R}$, with the requirement that the solution to the classical Riemann problem with flux $f_{l}$ and states $(\gamma,u_{r})$ only displays waves of negative speeds. We list below the different structure of the solution.
@@
The solution is the gluing of shock/rarefaction - SNS - rarefaction, or rarefaction - SNS - shock/rarefaction.
@@
The solution is the gluing of shock - SNS - constant $u_{r}$, or constant $u_{l}$ SNS - shock.
@@
The solution is the gluing of shock - SNS - rarefaction, or constant $u_{l}$ SNS - shock/rarefaction.

### v3

@@ 3 Proofs
To fix the ideas, the proofs are written in the convex case of Assumption (CVX).
@@ 3.1 Discontinuous Flux
We recall and adapt for our use results regarding conservation laws with discontinuous flux, borrowing elements from [AKR2011]. For the purpose of this section, let us fix $f_{l},f_{r}\in\mathbf{C}^{1}(\mathbb{R},\mathbb{R})$ two convex functions satisfying (CVX). Consider:
@@
Remark 3.1.
This structure will be applied in the scheme at each interface with $f_{l}(u)=H(x_{j},u)$ and $f_{r}(u)=H(x_{j+1},u)$.
Since the works of [AG2003, AJG2004, AGM2005, AKR2011, Diehl1995, GR1992, KRT2003, KT2004, Towers2000], it is now known that additionally to the conservation of mass (Rankine-Hugoniot condition)
$\text{for a.e.}\;t>0,\quad f_{l}(u(t,0-))=f_{r}(u(t,0+)),$
an entropy criterion must be imposed at the interface $\{x=0\}$ to select one solution. This choice is often guided by physical consideration. Among the most common criteria, we can cite: the minimal jump condition [GR1992, KMRWellBalanced, KMRTriangular], the vanishing viscosity criterion [BGS2019], or the flux maximization [AGM2005].
We call $\Phi_{l}$, resp. $\Phi_{r}$, the Kruzhkov entropy flux [Kruzhkov1970] associated with $f_{l}$, resp. with $f_{r}$, meaning that
$\forall a,b\in\mathbb{R},\quad\Phi_{l}(a,b)\coloneq\mathop{\rm sgn}(a-b)(f_{l}(a)-f_{l}(b)).$
In this section, define $\Phi=\Phi(x,u)$ as the Kruzhkov entropy flux associated with $F$:
@@
For the study of (3.1), we follow [AKR2011], where the traces of the solution at the interface $\{x=0\}$ are explicitly treated. It will be useful to have a name for the critical points of $f_{l,r}$, say $\alpha_{l,r}$. Notice that for all $y\in\mathopen{]}\min f_{r},+\infty\mathclose{[}$, the equation $f_{r}(u)=y$ admits exactly two solutions, $S_{r}^{-}(y)<\alpha_{r}<S_{r}^{+}(y)$. When $y=\min f_{r}$, then $S_{r}^{-}(y)=S_{r}^{+}(y)=\alpha_{r}$. This motivates the following definition.
@@
Define $\mathcal{G}\subset\mathbb{R}^{2}$ as the union of the following subsets:
$\mathcal{G}_{1}$ $\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}\geq\alpha_{l},\;k_{r}=S_{r}^{+}(f_{l}(k_{l}))\right\}$ (3.2)
$\mathcal{G}_{2}$ $\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}\leq\alpha_{l},\;k_{r}=S_{r}^{-}(f_{l}(k_{l}))\right\}$
$\mathcal{G}_{3}$ $\coloneq\left\{(k_{l},k_{r})\in\mathbb{R}^{2}\;:\;k_{l}>\alpha_{l},\;k_{r}=S_{r}^{-}(f_{l}(k_{l}))\right\}.$
We will show that $\mathcal{G}$ is a germ. The term was introduced in [AKR2011] and is used to describe an abstract subset having certain properties. The germ contains all the possible traces along $\{x=0\}$ of the solutions to (3.1). Notice that by construction, any couple in the germ satisfies the Rankine-Hugoniot condition. Conversely, some couples verifying the Rankine-Hugoniot condition have been excluded from the germ, more precisely the ones belonging to
@@
(ii) Dissipative inequality: for all $(u_{l},u_{r})\in\mathcal{G}$ and $(k_{l},k_{r})\in\mathcal{G}$, it holds:
$\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0.$ (3.3)
@@
$\bigl(\forall(k_{l},k_{r})\in\mathcal{G},\;\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0\bigr)\implies(u_{l},u_{r})\in\mathcal{G}.$ (3.4)
Proof.(i) Immediate from the construction.
(ii) Follows from a case by case study. Let us deal with one case. Assume for instance that $(u_{l},u_{r})\in\mathcal{G}_{1}$ and $(k_{l},k_{r})\in\mathcal{G}_{3}$. Recall that $u_{r}\geq\alpha_{r}\geq k_{r}$ and $u_{l},k_{l}\geq\alpha_{l}$. If $u_{l}\geq k_{l}$, then $\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})=0$. Otherwise, using the Rankine-Hugoniot condition,
$\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})=2(f_{l}(k_{l})-f_{l}(u_{l}))>0,$
where we used the monotonicity of $f_{l}$. The proof for the other cases follows the same idea.
(iii) We prove it by way of contradiction: let $(u_{l},u_{r})\in\mathbb{R}^{2}$ such that $f_{l}(u_{l})=f_{r}(u_{r})$ and assume that
@@
From the Rankine-Hugoniot condition, we deduce $u_{r}=S_{r}^{+}(f_{l}(u_{l}))$, with $u_{l}<\alpha_{l}$. Take $k_{r}\coloneq S_{r}^{-}(f_{l}(k_{l}))$ with $k_{l}\in\mathopen{]}u_{l},\alpha_{l}\mathclose{[}$. Clearly, $(k_{l},k_{r})\in\mathcal{G}_{2}\subset\mathcal{G}$, therefore by assumption, $\Phi_{l}(u_{l},k_{l})-\Phi_{r}(u_{r},k_{r})\geq 0$. But the choice of $k_{l}$ raises the contradiction:
@@
The theory of [AKR2011] proposes an abstract framework for the study of (3.1). On the other hand, the authors of [AGM2005, AJG2004] give a formula for the flux at the interface for Riemann problems of (3.1). In the following proposition, we link the two points of view. Let us adopt the notations:
@@
$\mathcal{R}(u_{l},u_{r})\coloneq{\left|F_{\mathrm{int}}(u_{l},u_{r})-f_{l}(u_{l})\right|}+{\left|F_{\mathrm{int}}(u_{l},u_{r})-f_{r}(u_{r})\right|}.$ (3.7)
@@
(ii) Set $q\coloneq\Phi_{r}(u_{r},k_{r})-\Phi_{l}(u_{l},k_{l})$. Keep in mind that $f_{l,r}$ is decreasing on $\mathopen{]}-\infty,\alpha_{l,r}]$ and increasing on $\alpha_{l,r},+\infty\mathclose{[}$. We proceed with a case by case study.
@@
Case 3: $u_{l}>k_{l}$ and $u_{r}\leq k_{r}$. Then $q=(f_{r}(k_{r})-f_{r}(u_{r}))+(f_{l}(k_{l})-f_{l}(u_{l}))$. We proceed by a case by case study, depending on which part of the germ $(k_{l},k_{r})$ belongs to. Notice that here, we have $\mathcal{R}(u_{l},u_{r})={\left|f_{l}(u_{l})-f_{r}(u_{r})\right|}$.
3.(a) $(k_{l},k_{r})\in\mathcal{G}_{1}$. Since $u_{l}>k_{l}\geq\alpha_{l}$, we write, using the monotonicity and $f_{l}(k_{l})=f_{r}(k_{r})$:
$q\leq f_{r}(k_{r})-f_{r}(u_{r})=(f_{l}(u_{l})-f_{r}(u_{r}))+(f_{l}(k_{l})-f_{l}(u_{l}))\leq f_{l}(u_{l})-f_{r}(u_{r})\leq\mathcal{R}(u_{l},u_{r}).$
3.(b) $(k_{l},k_{r})\in\mathcal{G}_{2}$. This time, $u_{r}\leq k_{r}\leq\alpha_{r}$, therefore:
$q\leq f_{l}(k_{l})-f_{l}(u_{l})=(f_{r}(u_{r})-f_{l}(u_{l}))+(f_{r}(k_{r})-f_{r}(u_{r}))\leq f_{r}(u_{r})-f_{l}(u_{l})\leq\mathcal{R}(u_{l},u_{r}).$
3.(c) $(k_{l},k_{r})\in\mathcal{G}_{3}$. Clearly, $q\leq 0$.
Case 4: $u_{l}\leq k_{l}$ and $u_{r}>k_{r}$. Similar to Case 3, so the details are omitted.   $\square$
@@
$\mathcal{R}(k,k)=\left\{\begin{aligned} {\left|\max\{\min f_{l},\min f_{r}\}-f_{l}(k)\right|}+{\left|\max\{\min f_{l},\min f_{r}\}-f_{r}(k)\right|}&\quad\text{if}\;\alpha_{r}\leq k<\alpha_{l}\\ {\left|f_{l}(k)-f_{r}(k)\right|}&\quad\text{otherwise.}\end{aligned}\right$.
We are now in position to properly give the definition of solutions for (3.1). The first definition is a Kruzhkov type entropy inequality where the solution is tested not against constant data, but against piecewise constant data. The second definition relies on the treatment of the traces of the solution along $\{x=0\}$. More precisely, for $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R})$, the strong traces of $u$ along $\{x=0\}$, if they exist, are the functions $\gamma_{L}u,\gamma_{R}u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[},\mathbb{R})$ such that
$\underset{r\to 0+}{\mathrm{ess\;lim}}\;\frac{1}{r}\int_{-r}^{0}{\left|u(t,x)-\gamma_{L}u(t)\right|}\mathinner{\mathrm{d}{x}}=0,\quad\underset{r\to 0+}{\mathrm{ess\;lim}}\;\frac{1}{r}\int_{0}^{r}{\left|u(t,x)-\gamma_{R}u(t)\right|}\mathinner{\mathrm{d}{x}}=0.$
@@
Define $\mathcal{G}$ as (3.2) and $\mathcal{R}$ as (3.7). Let $u_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$ and $u\in\mathbf{L}^{\infty}(\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R},\mathbb{R})$. Then the statements $\mathbf{(A)}$ and $\mathbf{(B)}$ below are equivalent.
$\mathbf{(A)}$ For all $\varphi\in\mathbf{C}_{\mathbf{c}}^{\infty}(\mathbb{R}^{+}\times\mathbb{R},\mathbb{R}^{+})$ and for all $(k_{l},k_{r})\in\mathbb{R}^{2}$, with $\kappa\coloneq k_{l}\mathbf{1}_{\mathopen{]}-\infty,0\mathclose{[}}+k_{r}\mathbf{1}_{\mathopen{]}0,+\infty\mathclose{[}}$, it holds:
$\int_{0}^{+\infty}\int_{\mathbb{R}}{\left|u-\kappa(x)\right|}\partial_{t}\varphi+\Phi(x,u,\kappa(x))\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}{\left|u_{o}(x)-\kappa(x)\right|}\varphi(0,x)\mathinner{\mathrm{d}{x}}$
@@
$\mathbf{(B)}$ The two following points are verified.
@@
$\int_{0}^{+\infty}\int_{\mathbb{R}}{\left|u-\kappa(x)\right|}\partial_{t}\varphi+\Phi(x,u,\kappa(x))\partial_{x}\varphi\;\mathinner{\mathrm{d}{x}}\mathinner{\mathrm{d}{t}}+\int_{\mathbb{R}}{\left|u_{o}(x)-\kappa(x)\right|}\varphi(0,x)\mathinner{\mathrm{d}{x}}\geq 0.$ (3.9)
(ii) For a.e. $t\in\mathopen{]}0,+\infty\mathclose{[}$, $(\gamma_{L}u(t),\gamma_{R}u(t))\in\mathcal{G}$.
When $\mathbf{(A)}$ or $\mathbf{(B)}$ holds, we say that $u$ is an entropy solution to (3.1) with initial datum $u_{o}$.
Proof.  See the proof of [AKR2011, Theorem 3.18]. It is worth mentioning that either (3.4) or (3.9) implies that $u$ admit strong traces along $\{x=0\}$. Indeed, since $f_{l}$ and $f_{r}$ are genuinely nonlinear in the sense that
$\forall s\in\mathbb{R},\quad\text{meas}\bigl(\left\{p\in\mathbb{R}\;:\;f_{l,r}^{\prime}(p)=s\right\}\bigr)=0,$
existence of strong traces is obtained from the works of [Vasseur2001, NPS2018].   $\square$
@@
    Let $u_{o},v_{o}\in\mathbf{L}^{\infty}(\mathbb{R},\mathbb{R})$. We denote by $u$, resp. $v$, the entropy solution to (3.1) with initial datum $u_{o}$, resp. $v_{o}$. Then, there exists $L>0$ such that for all $R>0$ and for all $t>0$,
    $\int_{{\left|x\right|}\leq R}{\left|u(t,x)-v(t,x)\right|}\mathinner{\mathrm{d}{x}}$ $\leq\int_{{\left|x\right|}\leq R+Lt}{\left|u_{o}(x)-v_{o}(x)\right|}\mathinner{\mathrm{d}{x}}$ (3.10)
    $\int_{{\left|x\right|}\leq R}(u(t,x)-v(t,x))^{+}\mathinner{\mathrm{d}{x}}$ $\leq\int_{{\left|x\right|}\leq R+Lt}(u_{o}(x)-v_{o}(x))^{+}\mathinner{\mathrm{d}{x}}.$
Proof.  The proofs of the stability bounds can be found in [AKR2011, Theorems 3.11-3.19]. Regarding the existence, one can build a simple finite volume scheme and prove its convergence, see for instance [AJG2004]. The flux to put at the interface $\{x=0\}$ in the scheme is precisely $F_{\mathrm{int}}$.   $\square$
@@
$F_{\mathrm{int}}(u_{l},u_{r})=\left\{\begin{array}[]{ccl}\max\{\min f_{l},f_{r}(u_{r})\}&\text{if}&u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\quad\text{(I)}\\ \max\{\min f_{l},\min f_{r}\}&\text{if}&u_{l}\leq\alpha_{l}\;\text{and}\;u_{r}>\alpha_{r}\quad\text{(II)}\\ \max\{f_{l}(u_{l}),f_{r}(u_{r})\}&\text{if}&u_{l}>\alpha_{l}\;\text{and}\;u_{r}\leq\alpha_{r}\quad\text{(III)}\\ \max\{f_{l}(u_{l}),\min f_{r}\}&\text{if}&u_{l}>\alpha_{l}\;\text{and}\;u_{r}>\alpha_{r}\quad\text{(IV)}.\end{array}\right$. (3.11)
The value of $F_{\mathrm{int}}(u_{l},u_{r})$ imposes the value of one of the traces at the interface. For instance, if $F_{\mathrm{int}}(u_{l},u_{r})=f_{r}(u_{r})$, then the right-hand trace of the solution along $\{x=0\}$ is equal to $\gamma_{r}\coloneq u_{r}$. To compute the left-hand trace, one solves $f_{l}(\gamma)=f_{r}(\gamma_{r})$, $\gamma\in\mathbb{R}$. This equation admits two solutions, $\gamma_{l}^{-}\leq\alpha_{l}\leq\gamma_{l}^{+}$, but only one of them is such that the classical Riemann problem with flux $f_{l}$ and states $(\gamma_{l},u_{r})$ only displays waves of negative speeds. This determines the left-hand trace. Finally, the solution to the Riemann problem is obtained by patching together:
  • the solution to the classical Riemann problem with flux $f_{l}$ and states $(u_{l},\gamma_{l})$ on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}^{-}$ ;
  • the stationary (non entropic) shock connecting $(\gamma_{l},\gamma_{r})$ at $x=0$ ;
  • the solution to the classical Riemann problem with flux $f_{r}$ and states $(\gamma_{r},u_{r})$ on $\mathopen{]}0,+\infty\mathclose{[}\times\mathbb{R}^{+}$ .
As an illustration, consider the data
$f_{l}(u)\coloneq\frac{u^{2}}{2},\quad f_{r}(u)\coloneq u^{2},\quad u_{o}^{(1)}\equiv-1,\quad u_{o}^{(2)}(x)\coloneq\left\{\begin{array}[]{ccl}-1&\text{if}&x<0\\ 1&\text{if}&x>0.\end{array}\right$.
The values of $u_{o}^{(1)}$ lead to Case (I). The interface flux is $f_{r}(-1)$, therefore the right-hand trace of the solution is $\gamma_{r}=-1$. Then,
$f_{l}(\gamma)=f_{r}(\gamma_{r})\iff\gamma\in\{-\sqrt{2},\sqrt{2}\}.$
The left-hand trace of the solution is $\gamma_{l}\coloneq-\sqrt{2}$ because the solution to the Riemann problem with flux $f_{l}$ and states $(-1,\sqrt{2})$ is a rarefaction wave presenting a positive speed. Consequently, the solution is
$u^{(1)}(t,x)=\left\{\begin{array}[]{ccl}-1&\text{if}&x<-\frac{t}{2(\sqrt{2}-1)}t\\ -\sqrt{2}&\text{if}&-\frac{t}{2(\sqrt{2}-1)}t<x<0\\ -1&\text{if}&x>0.\end{array}\right$.
On the other hand, the values of $u_{o}^{(2)}$ lead to Case (II). The interface flux is $0=f_{l}(0)=f_{r}(0)$. We deduce that both traces of the solution at the interface are equal to 0. Therefore, the solution is obtained by patching together two rarefaction waves:
$u^{(2)}(t,x)=\left\{\begin{array}[]{ccl}-1&\text{if}&x<-t\\ \frac{x}{t}&\text{if}&-t<x<0\\ \frac{x}{2t}&\text{if}&0<x<2t\\ 1&\text{if}&x>2t.\end{array}\right$.
Below are listed the possible structures of the solution to a Riemann problem with discontinuous flux.
@@
The solution is the gluing of shock/rarefaction - SNS - rarefaction or rarefaction - SNS - shock/rarefaction.
@@
The solution is the gluing of shock - SNS - constant $u_{r}$, or constant $u_{l}$ - SNS - shock.
@@
The solution is the gluing of shock-SNS-rarefaction, or constant $u_{l}$ - SNS - shock/rarefaction.